40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Fri Jul 9 12:25:25 2010, kiwamu, Update, Green Locking, SHG on PSL table PSL_doubling.png
    Reply  Tue Jul 13 11:00:29 2010, kiwamu, Update, Green Locking, SHG on PSL table : optimum temeprature power_temp.png
       Reply  Tue Jul 13 11:20:07 2010, Dmass, Update, Green Locking, SHG on PSL table : optimum temeprature 
Message ID: 3203     Entry time: Tue Jul 13 11:00:29 2010     In reply to: 3188     Reply to this: 3204
Author: kiwamu 
Type: Update 
Category: Green Locking 
Subject: SHG on PSL table : optimum temeprature 

The optimum temperature for the doubling crystal on the PSL table was found to be 36.8 deg.

I scanned the temperature of the crystal from 44 deg to 29 deg, in order to find the optimum temperature where the frequency doubled power is maximized. 

 

(method) 

The method I performed is essentially the same as that Koji did before (see this entry).

(1) First of all, I enabled the PID control on the temperature controller TC200 and set the temperature to 44 deg.

(2) After it got 44 deg, I disabled the PID control.

(3) Due to the passive cooling of the oven, the temperature gradually and slowly decreased. So it automatically scans the temperature down to the room temperature.

(4) I recorded the power readout of the power meter: New Port 840 together with the temperature readout of TC200. The power meter was surely configured for 532 nm.

 

(result)

The measured data are shown in the attachment. 

The peak was found at T=36.8 deg where the power readout of  532 nm was 605 uW.

Compared with Koji's past data (see this entry), there are no big side lobes in this data. I am not sure about the reason, but the side lobes are not critical for our operation of the green locking.

 

 (to be done)

 Adjustment of the PID parameters

Attachment 1: power_temp.png  53 kB  | Show | Show all
ELOG V3.1.3-